OpenStudy (anonymous):

Which expression shows the value of cosec x°? Answer adjacent ÷ opposite adjacent ÷ hypotenuse hypotenuse ÷ opposite hypotenuse ÷ adjacent

jimthompson5910 (jim_thompson5910):

sin(x) = opposite/hypotenuse 1/sin(x) = 1/(opposite/hypotenuse) csc(x) = hypotenuse/opposite

OpenStudy (anonymous):

coescx = H/P or opposite/hypotenuse